let a, b and c are unequal real positive numbers such that 2b = a + c, then roots of ax^2 + 2bx + c = 0 are


(i) real and equal (ii) Real and distinct (iii) imaginary (iv) nothing definite can be said


Plz explain

Grade:10

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

2b=a+c

Sum of the roots=-2b/a

Product=c/a

This shows the roots are real and distinct...

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com

vikas askiitian expert
13 years ago

for any quadratic eq  ax2+bx+c roots are given by

 X = {-b+_ D)/2a

D =discriminant = (b2-4ac)1/2

our eq is ax2+2bx+c so

D = {(2b)2-4ac)1/2 ........1

2b = a+c  (given)

so , (2b)2 = (a+c)2 ...........2

putting  eq2 in eq1

D = { (a+c)2 - 4ac}1/2  =  { a2+c2-2ac}1/2 ={(a-c)2}1/2 =a-c

now roots are

 X = (-2b +_ D}/2a = {-2b + (a-c)}/2a     or        {-2b-(a-c)}/2a                                   

                           = -c/a                      or        -1                                  (on putting 2b =a+c )

so the roots are unequal & real

option (ii) is correct

Fawz Naim
13 years ago

the equation is ax^2+2bx+c=0

the discriminant is= B^2-4AC

A=a,B=2b,C=c

discriminant= 4b^2-4ac

but 2b=a+c, therefore (2b)^2=(a+c)^2

= (a+c)^2-4ac

=a^2+c^2+2ac-4ac

=a^2+c^2-2ac

=(a-c)^2, which is always greater than zero

therefore roots are real and distinct